OpenStudy (anonymous):

a) find the value or values of x that make the denominator 0 b) keeping the restrictions in mind, solve the equation (1/x-4) - (5/x+2) = (6/x^2+x-6)

OpenStudy (anonymous):

are all the parenthisis on their correct place?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

http://i.imgur.com/CxmjZ.png Is this your expression?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

parenthesis means its all one expression inside

OpenStudy (anonymous):

then you dont have the parenthisis on the correct place. Cant help you again

OpenStudy (anonymous):

so if i say (1/x-4) it looks like this 

OpenStudy (anonymous):

if i meant what you said it would read (1/x)+4

OpenStudy (anonymous):

why dont you just type like this like a normal person: 1/(x-4) - 5/(x+2) = 6/(x^2+x-6)

OpenStudy (anonymous):

you know what the problem is now can you help?

OpenStudy (anonymous):

im on it. just a sec

OpenStudy (anonymous):

a) x=-3 x=-2 x=2 x= 4 b) http://i.imgur.com/16Ote.png

OpenStudy (anonymous):

the books given answers were a) -2,4 and b) 0

OpenStudy (anonymous):

i understand the -2,4 but not the 0

OpenStudy (anonymous):

in which grade are you?

OpenStudy (anonymous):

sophomore college

OpenStudy (anonymous):

do you already know what imaginary numbers are?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hmm.. your answers are for the left side of the equation, right? cause if X = -3 or x= 2 on the right side, the dominator is also 0

OpenStudy (anonymous):

oh sorry the book answer is no solution, not zero

OpenStudy (anonymous):

no -3 works for the right side

OpenStudy (anonymous):

9-6-8 isnt 0

OpenStudy (anonymous):

9+6-8*

OpenStudy (anonymous):

(-3)^2 -3-6 9 -9 = 0 -3 works

OpenStudy (anonymous):

i fear ive wasted your time, i accidentally entered the wrong numbers for the last 3 numbers on the right side looking at another problem

OpenStudy (anonymous):

the right side is actualy 6/(x^2-2x-8)

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

ohhh so that only changes.... EVERYTHING a) x= 4 x= -2 b) no solution

OpenStudy (anonymous):

hahahha

OpenStudy (anonymous):

feel free to ignore may questions in the future

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

can you explain how its no solution though?

OpenStudy (anonymous):

as you can see from the plot, they never intersect each other http://i.imgur.com/B5ubz.png

OpenStudy (anonymous):

well how can you find that without graphing nine lines?

OpenStudy (anonymous):

try to solve for X and you'll see you wont get anywhere

OpenStudy (anonymous):

i ended up with 4

OpenStudy (anonymous):

ohhh but it cant be 4!

OpenStudy (anonymous):

so its undefined.. got it

OpenStudy (anonymous):

gottcha